The rate of decomposition of azomethane is studied by monitoring the partial pressure of the reactant as a function of time. The data obtained at 300°C are listed on the right.

Determine the rate constant of the reaction at this temperature.

time:

0, 100, 150,200,250,300

P (mmHg):

284,220,193, 170,150,132
Answer #1

P vs T plot is

o 100 150 200 250 300 284 220 193 170 150 132 5 y = 283.7586e-0.0026x R2 = 1.0000 7T 0 50 100 150 200 250 300 350 17The plot follows exponential curve with R2 =1

then it follows first order kinetics

now

lnP vs T ( time) plot is

0 100 150 200 5.65 5.39 5.26 5.14 5.01 4.88 250 300 In P y = -0.0026x + 5.6479 R2 = 0.9999 0 50 100 150 200 250 300 350 1/T

Bestline equation is

y = - 0.0026 x + 5.6479 (1)

Integrated rate law of first order kinetics is

ln [A] = - kt + ln[A]0   (2)

so comparing Eq. 1 and Eq. 2

Rate constant (K) = 0.0026 s-1

[

note : unit of rate constant is time -1 . no unit is given in the question so commonly used unit s-1 is given. ]